An object falls from rest, and s =-16t² where s feet is the distance of the object from the starting point at t seconds, and the positive direction is upward. If a stone is dropped from a building 256 ft high, find


A. the instantaneous velocity of the tone 1 sec after it is dropped;


B. the instantaneous velocity of the stone 2 sec after it is dropped;


C. how long it takes the stone to reach the ground;


D. the instantaneous velocity ot the stone when it reaches the ground.


1
Expert's answer
2021-03-11T06:12:18-0500

Given,

"s= -16t^2\\"

"g= 32ft\/sec^2"

"S= 256ft"


(a) Instantaneous velocity of the stone= "g\times t" ft/sec

So, at t=1sec, the instantaneous velocity of stone = "g\times t=32\times 1=32ft/sec"


(b) Similarly, at t=2sec, the instantaneous velocity of stone ="g\times t_{2sec}=32\times 2= 64 ft/sec"


(c) Total distance travelled to reach the ground "S=\dfrac{1}{2}gt^2"

So, total time it takes to reach ground "t=\sqrt{\dfrac{2S}{g}}=\sqrt{\dfrac{2\times256}{32}}=\sqrt{16}=4 sec"


(d) The stone reaches the ground at t=4sec

So, instantaneous velocity of stone when it reaches ground


"v= g\times t_{4sec}=32\times 4=128ft/sec"